Okay here is my question...
The sum of Jamie's age and Devon's age is 30. Ten years ago the product of their ages was 24. If Jamie is younger than Devon, what are their ages?

Answer by stanbon(75887)

The sum of Jamie's age and Devon's age is 30. Ten years ago the product of their ages was 24. If Jamie is younger than Devon, what are their ages?
-------------------
NOW DATA:
Let Jamie's age be "x"
Devon's age is "30-x"
--------------
Ten Years ago DATA:
Jameie's age was "x-10"
Devon's age was "30-x-10" = "20-x"
EQUATION:
Ten yrs. ago the product of their ages was 24
(x-10)(20-x)=24
20x-x^2-200+10x=24
-x^2+30x-200=24
x^2-30x+224=0
(x-16)(x-14)=0
x=16 or x=14
x=[30+-sqrt(900-4(224)]/2
x=[30+-2]/2
x=16 or x=14
Jamie is younger than Devon so Jamie is 14.
Cheers,
Stan H.